OpenStudy (anonymous):

Write the given sentence as an equation. Tim’s age in 7 years will be three times what it was 19 years ago. A. 3(t + 7) = t – 19 B. t + 19 = 3(t – 7) C. t + 7 = 3(t – 19) D. 3(t + 19) = t – 7

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

I think it is C

OpenStudy (anonymous):

It's B! Why, C is incorrect.. example: lets say he's 20 now.. t=20 t + 7= 3(t-19) 20+ 7 = 3(20-19) 27= 3(1) 27= 3, which isnt correct.. So, i dont think its C Its B example: t=20 t + 19 = 3(t – 7) 20+19 = 3(20-7) 39 = 3(13) 39 = 39, which is correct!